OpenStudy (anonymous):

***Medal*** The total snowfall in Buffalo, NY one year was one hundred twenty inches. If the total snowfall the following year was ninety inches, what was the percent change? Show and explain your solution.

OpenStudy (anonymous):

120-90 / 100% = ?

OpenStudy (anonymous):

lets see lets put it as 120 was 100%

OpenStudy (anonymous):

Its pretty simple :3 but i think instead of dividing its multiplying not sure

OpenStudy (anonymous):

If you know the answer it would be very helpful.. I got really behind in all my classes and im just trying to get everything done in 2 weeks

OpenStudy (mathmate):

Percentage change = difference/old value * 100% If last year's rainfall was 12 cm, and this year is 10 cm, then the percentage change = difference/old * 100% =(new-old)/old * 100% = (10-12)/12 *100% =-2/12 *100% = -16.67%

OpenStudy (mathmate):

@Football_babe2800 Just to make sure, the answer I gave (-16.67%) is for the example I gave. You can follow the same steps to get the answer to the problem posted. :)